Below is an incomplete contribution margin income statement for Barry's Coffee Cakes. Use this information to answer the following questions.

BARRY'S COFFEE CAKES
Contribution Margin Income Statement
20,000 units 25,000 units
Total Total Per Unit Percents of Sales
Sales revenue $ 150,000
Less: Variable costs 50,000
Contribution margin 100,000
Less: Fixed costs 30,000
Net operating income $ 70,000

Knowledge Check 01

What is the unit contribution margin?

  • $7.50

  • $2.50

  • $5.00

  • $3.00

Knowledge Check 02

What is the contribution margin ratio?

  • 33%

  • 50%

  • 40%

  • 67%

Knowledge Check 03

If there is no change in fixed costs, what will the net operating income be when 25,000 units are sold?

  • $95,000

  • $125,000

  • $70,000

  • $87,500

Below is the data from a regression analysis performed by Earth Right Spa on its overhead costs and clients for the past year. Use this information to answer the following questions.

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.949
R Square 0.901
Adjusted R Square 0.891
Standard Error 1102.512
Observations 12.000
ANOVA
df ss MS f Significance F
Regression 1 111011767.37 111011767.37 91.33 0.00
Residual 10 12155332.63 1215533.26
Total 11 123167100.00
Ceofficients Standard Error tStar P-value Lower 95% Upper 95% Lower 95.0% Upper 95%
Intercept 6825.84 724.43 9.42 0.00 5211.70 8439.98 5211.70 8439.98
Guests (X) 32.86 3.44 9.56 0.00 25.20 40.52 25.20 40.52

Knowledge Check 01

Identify the fixed overhead cost per month from the data provided.

  • $8,364.72

  • $946.56

  • $6,825.84

  • $6,546.54

Knowledge Check 02

What is the variable overhead cost per client served?

  • $724.43

  • $25.20

  • $40.52

  • $32.86

Knowledge Check 03

What will be the total overhead cost if 100 clients are served?

  • $10,111.84

  • $79,269.07

  • $9,345.86

  • $10,878.24

Month Customers Total Overhead Cost
March 5,000 $ 85,000
April 7,500 $ 91,500
May 8,000 $ 94,000
June 9,000 $ 97,000
July 6,000 $ 87,800
August 6,500 $ 92,700

Knowledge Check 01

What is the total overhead cost in the month with the highest activity?

  • $85,000

  • $91,500

  • $97,000

  • $94,000

Knowledge Check 02

What is the variable cost per unit?

  • $2.70

  • $4.00

  • $3.50

  • $3.00

Knowledge Check 03

What is the total fixed overhead cost?

  • $85,000

  • $80,000

  • $50,000

  • $70,000

The span at which the cost behaviors are expected to hold true is called:
Answer #1

Unit contribution margin = Total Contribution margin/units

= 100,000/20,000

= 5.00

Contribution margin ratio = Total Contribution margin/Total sales revenue

= 100,000/150,000

= 67%

Net operating income = Total Contribution margin - Fixed cost

= (5*25,000) - 30,000

= 95,000